ple solve this qutn

prove that the locus of center of the circle



1\2(x2+y2)+ x cosA + y sinA - 4 = 0


is x2+ y2 = 1





Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear rohan

1\2(x2+y2)+ x cosA + y sinA - 4 = 0

or (x2+y2)+ 2x cosA + 2y sinA - 8 = 0

center of circle is (-cosA ,-sinA)

let center of circle is (h,k)

 so h=-cosA   and k=-sinA

h2+k2 =sin2A +cos2A

h2+k2 =1

so locus

x2+y2 =1


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin